A chemist has two solutions of hydrochloric acid in stock: a 50% solution and an 80% solution. How much of each should be used to obtain 100 milliliters of a 68% solution?
Answer by ikleyn(52867)

A chemist has two solutions of hydrochloric acid in stock: a 50% solution and an 80% solution. How much of each should be used to obtain 100 milliliters of a 68% solution?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

   x +    y = 100,         (1)     (total mass equation)
0.5x + 0.8y = 0.68*100.    (2)     (hydrochloric acid content equation)

or

   x +    y = 100,         (1')
0.5x + 0.8y = 68.          (2')

Express y = 100 - x from (1') and substitute it into (2').

Then solve the obtained equation for x.

After finding x, substitute it into (1) and solve for y.